# Phase separation of chemokinetic active particles

### v1

@@ III.1 BMR facilitates phase separation
It is notable that the phase boundary changes discontinuously from $D_{\text{vanilla}}$ to $D_{\text{BMR}}$ as the chemical consumption is turned on. When the chemicals are irrelevant, MIPS occurs only via the deceleration of particles with increasing density. In contrast, when $\lambda>0$, high local particle density and low local chemical concentration can support each other via positive feedback, creating density and concentration modulations against diffusive currents. When the particle-chemical coupling $\lambda$ is small, the density-concentration feedback can overcome only small diffusion currents, which means that only long-wavelength modulations can be maintained. Nonetheless, whenever $\lambda>0$, there exists a finite wavelength above which the density and concentration modulations can be amplified against diffusive currents. Thus, the mechanism of MIPS for $\lambda>0$ is fundamentally different from the corresponding mechanism for $\lambda=0$, which explains the discontinuous change of the phase boundary.
Figure 3: Phase diagram of the hydrodynamic model in the AMR at $n_{0}=20$ and $\zeta=22$. The LSA predictions for the phase boundaries are indicated by the dotted (H–S boundary), the dash-dotted (H–O boundary), and the dashed (S–O boundary) lines. For comparison, snapshots of particle distribution (with darker regions indicating higher density) obtained by hydrodynamic simulations are shown together. The values of $\lambda$ and $D_{\mathrm{c}}$ for each snapshot correspond to the coordinates of its center.
@@
To check the validity of the LSA, we perform numerical simulations of Eq. (II) via the pseudo-spectral method. In Fig. 1(b), we show the phase boundary (indicated by a dotted curve) predicted by the LSA together with snapshots of the numerical simulations at $t=2\times 10^{3}$ as the global mean concentration $n_{0}$ and the deceleration parameter $\zeta$ are varied within the physical regime (the grayed-out area corresponds to the unphysical regime where $v(\rho,n)<0$). Note that the coordinates of the center of each snapshot indicate the associated values of $n_{0}$ and $\zeta$. Clearly, the predicted boundary correctly separates the homogeneous snapshot from the phase-separated ones.
@@
Figure 4: Phase-separated patterns of the hydrodynamic model in the AMR. Particle density $\rho(\mathbf{r},t)$ (blue) and the magnitude of local vorticity $|\boldsymbol{\omega}(\mathbf{r},t)|$ (green) are shown for (a) diffusive droplets ($\lambda=0.0025$, $D_{\mathrm{c}}=1.75$, $n_{0}=10$, $\zeta=22$), (b) ballistic bands ($\lambda=0.01$, $D_{\mathrm{c}}=1$, $n_{0}=17.5$, $\zeta=28.5$), and (c) radiating ripples ($\lambda=0.01$, $D_{\mathrm{c}}=1$, $n_{0}=23.5$, $\zeta=28.5$). (d, e) Mean cluster size in the S phase as $\lambda$ is varied. The error bars indicate the standard deviation of cluster size. (f) Mean vorticity as $D_{\mathrm{c}}$ is varied for $\lambda=0.0025$ and (inset) as $\lambda$ is varied for $D_{\mathrm{c}}=0.25$. Note that $n_{0}=20$ and $\zeta=22$ in (d–f).
We also show how the phase-separated patterns evolve with time in Fig. 2. For the vanilla MIPS, whether phase separation occurs via bubble or cluster formation depends on the global density $\rho_{0}$ of particles. However, by introducing chemokinetic effects, the particles can be made to form clusters [as shown in Fig. 2(a)] or bubbles [as shown in Fig. 2(b)] even at the same value of $\rho_{0}$ by adjusting the global chemical concentration $n_{0}$. This is thanks to the strong anticorrelation between $\rho(\mathbf{r},t)$ and $n(\mathbf{r},t)$, which makes the chemicals form clusters (bubbles) where the particles form bubbles (clusters). Since large (small) $n_{0}$ favors bubble (cluster) formation of the chemicals, the particles are driven to form clusters (bubbles). We also note that the length scale of the bubbles and clusters gradually increases in time, which suggests the coarsening dynamics driven by the Ostwald process.
III.2 AMR induces small, oscillating clusters
@@
The LSA for the AMR can be done similarly, but it differs from that for the BMR in that $\mathcal{B}$ and $\mathcal{C}$ in Eq. (8) may have different signs ($\mathcal{B}<0$ and $\mathcal{C}>0$). The eigenvalues $\omega_{\pm}$, obtained in Eq. (9), may then have nonzero imaginary parts. Thus, the AMR exhibits the O phase in addition to the H and the S phases, which leads to some extra complications in predicting the phase boundaries. Nonetheless, the LSA can be carried out exactly, as detailed in Appendix A. Through the analysis, we can prove that the presence of the particle-chemical coupling ($\lambda>0$) always broadens the H phase. This confirms our previous expectation (discussed in Sec. II) that the AMR suppresses the MIPS.
@@
Examples of the $\zeta$$n_{0}$ and the $D_{\mathrm{c}}$$\lambda$ phase diagrams for the AMR are shown in Fig. 1(c) and 3, respectively. The LSA prediction for the H–S phase boundary is marked by the dotted line, the H–O boundary by the dash-dotted line, and the S–O boundary by the dashed line. For comparison, we also show the snapshots of the particle density, which are obtained by numerical simulations of Eqs. (II) via the pseudo-spectral method. shown together for comparison, where the coordinates of the center of each square corresponds to the parameters used in the simulation. The predicted H–S and the H–O boundaries, for which we expect the LSA to be exact, are in good agreement with the numerics. In contrast, the O phase seems to be broader than indicated by the predicted S–O boundary, as exemplified by the lower left corner of Fig. 3. Since the LSA is valid only up to linear order in density and concentration fluctuations, it is only natural that the true S–O boundary would deviate from the LSA prediction.
@@
appearing in Eq. (II). We note that high vorticity tends to coincide with regions where patterns evolve more rapidly. With these in mind, we can identify at last three notable patterns, namely diffusive droplets [see Fig. 4(a)] well within the S phase, ballistic bands near the boundaries of the O phase [see Fig. 4(b)], and radiating ripples well within the O phase [see Fig. 4(c)]. It seems that the latter two patterns are essentially the same, the only difference being that radiating ripples have a smaller length scale. This is also corroborated by the LSA, which shows that the unstable oscillating fluctuations tend to exhibit shorter wavelengths in the regimes of radiating ripples compared to ballistic bands.
@@
Let us first look into the properties of the diffusive droplets. As indicated by Fig. 4(d), the average number of particles forming each droplet (which we call the mean cluster size) decreases with $\lambda$ and increases with $D_{\mathrm{c}}$. The former is in agreement with our intuition that the particle-chemical coupling in the AMR suppresses the MIPS. The latter is due to the high $D_{\mathrm{c}}$ strongly suppressing the short-wavelength undulations of the chemical concentration. Since all patterns are formed by the particle density and the chemical concentration sharing the same characteristic length scale, this effect also drives the particles to form larger droplets. However, as implied by the snapshots in Fig. 3, the mean cluster size starts to increase again when $D_{\mathrm{c}}$ decreases below $2.5$. This is due to the droplets becoming increasingly motile and actively colliding with each other as the system approaches the S–O boundary.
All these behaviors are in stark contrast to the case of the BMR, where the MIPS eventually forms a single macroscopic particle cluster via coarsening, as exemplified by Fig. 2. Indeed, Fig. 4(e) indicates that the mean cluster size does not change appreciably with the system size $L$ when $\lambda>0$. In other words, the size of each droplet stays finite, and the S phase of the AMR exhibits microphase separation.
Now, we examine how the system changes quantitatively as it transitions from the S phase with diffusive droplet to the O phase with ballistic bands and radiating ripples. As already shown in Figs. 4(a)–(c), the latter two patterns exhibit spatially extended regions of high vorticity, reflecting their large-scale evolution in time. In contrast, the high-vorticity regions for the diffusive droplets are limited to the droplet boundaries. This implies that the mean vorticity
@@
can be used as an indicator of the O phase. Indeed, Fig. 4(f) shows that the mean vorticity dramatically increases as $D_{\mathrm{c}}$ is decreased for $\lambda=0.0025$, crossing the S–O boundary shown in Fig. 3. Notably, even before reaching the S–O boundary, the mean vorticity gradually increases as $D_{\mathrm{c}}$ is decreased, which is then followed by a sharp dip around $D_{\mathrm{c}}=2$. This nonmonotonic behavior might be due to the individual droplets becoming increasingly motile as $D_{\mathrm{c}}$ is reduced, which then merge together by collisions and shorten the boundaries, thus reducing the high-vorticity regions as well. The inset of Fig. 4(f), which shows the mean vorticity for $D_{\mathrm{c}}=0.25$ as $\lambda$ is varied, similarly captures the O phase lying between the S and the H phases in the lower part of Fig. 3.
Figure 5: Evolution of particle density $\rho(\mathbf{r},t)$ and chemical concentration $n(\mathbf{r},t)$ in the hydrodynamic simulation starting from the initial state with two particle clusters and the uniform concentration. The parameters are given by $n_{0}=20$, $\zeta=25$, $D_{\mathrm{c}}=0.1$, and $\lambda=0.5$, which corresponds to the S phase.
With the qualitative differences between different patterns thus established, we move onto understanding their underlying dynamics. We first examine how the microphase separation occurs in the S phase, starting from a configuration of two large particle clusters and the uniform chemical distribution. As depicted in Fig. 5, these large particle clusters quickly disintegrate into smaller clusters (including some ring-like structures), while the chemical develops a pattern that overlaps almost exactly with the particle distribution. Then these intermediate particle-chemical clusters gradually shrink, letting the other particle-chemical clusters grow at their expense until all clusters have roughly the same size. This suggests the reverse Ostwald process underlying the microphase separation, which was also found to play a crucial role in the microphase separation of generic single-species scalar active matter Tjhung et al. (2018).
@@
Figure 6: Profiles of particle density $\rho(\mathbf{r})$ and chemical concentration $n(\mathbf{r})$ in (a, b) the S phase ($D_{\mathrm{c}}=2.25,\lambda=0.0025,\rho_{0}=20,\zeta=22$) and (c, d) the O phase ($D_{\mathrm{c}}=1,\lambda=0.01,\rho_{0}=25.5,\zeta=17.5$). Note that (b) and (d) are cross sections of (a) and (c), respectively. The arrows in (c) and (d) indicate the directions in which the patterns move.
@@
We also look into how the particle-chemical coupling creates the traveling patterns in the O phase. When $D_{\mathrm{c}}$ is large, the chemical concentration is slaved to the particle density, so that their undulations overlap almost exactly as shown in Fig. 6(a) and (b). This leads to the diffusive droplets observed in the S phase. In contrast, when $D_{\mathrm{c}}$ is sufficiently small, the particle and the chemical dynamics may have comparable time scales, in which case the following scenario becomes possible. Suppose that the particle density undulations are slightly shifted to the right with respect to those of the chemical concentration. Then, for each density peak, the chemical concentration is lower (higher) on the right (left), decreasing (increasing) the speed of particles there. This means that particles tend to accumulate on the right and evaporate on the left, making the peak travel to the right. The concentration peak follows the density peak but cannot catch up for a prolonged period of time since $D_{\mathrm{c}}$ is small. This is precisely what happens for the density and the concentration fields shown in Figs. 6(c) and (d), which exhibit band-like patterns traveling to the right.
It is important to note that these traveling patterns in the O phase are consequences of the nonreciprocity reflected in the different signs of $\mathcal{B}$ and $\mathcal{C}$ appearing in Eq. (8). Known examples of traveling bands share the same property, such as the Vicsek model Vicsek et al. (1995) in which the interpaticle interactions are nonreciprocal, and the multispecies scalar active matter obeying the non-reciprocal Cahn–Hilliard dynamics Saha et al. (2020); Pisegna et al. . In a sense, our model can also be regarded as a multispecies scalar active matter with continuously varying self-propulsion speed. However, unlike the previous examples, the band patterns of our model are not necessarily straight or aligned in a single direction, as clearly exemplified in Figs. 4(b) and (c).
Figure 7: (a) Fraction of particles belonging to the largest cluster (identified using the DBSCAN algorithm) as the global chemical concentration $n_{0}$ is varied in the particle-based model with the WCA potential. Error bars are obtained using six different realizations. Snapshots of the particle (blue) and chemical (red) distributions are shown together. We use $\sigma=1/2$, $u_{0}=100$, $\lambda=0.1$, $D_{\mathrm{c}}=1$, and $L=200$, with the other parameters identical to those used in Fig. 1. (b–d) Histograms of local particle density as $n_{0}$ is varied for (b) the BMR, (c) the vanilla model, and (d) the AMR. All physical quantities and snapshots are measured at time $t=200$.
IV Particle-based simulations
Thus far, we have discussed the phase separation and pattern formation behaviors of the hydrodynamic model. It still remains to be checked whether the same behaviors persist for the original particle-based models from which the hydrodynamic model is derived. In this section, we address the issue using the ABPs interacting mechanically via the Weeks–Chandler–Andersen (WCA) potential and the quorum-sensing ABPs whose parameters closely match those of the hydrodynamic model.
IV.1 Mechanically interacting ABPs
Here, we directly simulate the particle-based model following Eqs. (II)–(4), with the short-range repulsion between particles given by the WCA potential
where $\sigma$ is the effective radius of each particle. We note that the Dirac delta function appearing in Eqs. (3) and (4) are discretized in the simulation, so that its value becomes $1$ only when the particle is in the grid corresponding to $\mathbf{r}$ and 00 otherwise.
To verify whether the BMR (AMR) facilitates (suppresses) the MIPS as shown for the hydrodynamic model, we let the above model evolve from a homogeneous initial state to $t=200$ and observe its phase separation behaviors. In Fig. 7(a), we show the fraction of particles belonging to the largest cluster (also called the giant cluster) for the BMR, the AMR, and the vanilla model using the same set of parameters (except for $\lambda=0$ used in the vanilla model). The BMR develops a larger particle cluster compared to the vanilla model, with the chemical concentration becoming higher outside the particle cluster. In contrast, the AMR strongly suppresses the growth of clusters, with the chemical concentration becoming higher inside the particle cluster. It is also notable that the giant clusters formed by the BMR have smooth interfaces, while those formed by the AMR tends to be of rugged shapes. This rules out the presence of the Ostwald process driven by the positive interfacial tension in the AMR, which is crucial for the large-scale phase separation in the BMR.
Using the histograms of the local particle density, we also check whether phase separation occurs for the BMR [see Fig. 7(b)], the vanilla model [see Fig. 7(c)], and the AMR [see Fig. 7(d)] at different values of $n_{0}$. For the BMR, the separation between dense and sparse regions occurs at the smallest value of $n_{0}$ among the three cases, and the density gap at a fixed value of $n_{0}$ is also the largest. For the AMR, on the contrary, the phase separation can hardly be detected by the histograms. In other words, the AMR does not exhibit phase separation, or the dense particle clusters are too small to be captured by the histograms.
@@ IV.1 Mechanically interacting ABPs
These results indicate that the BMR facilitates the large-scale MIPS, whereas the AMR suppresses the MIPS and allows only microphase separation. These results are consistent with the behaviors of the hydrodynamic model. However, the O phase of the AMR is not clearly observable in this particle-based model. This may be because the microscopic noises, which are neglected in the hydrodynamic model, overwhelm the oscillating patterns. This raises the question of whether the O phase of the AMR identified by the LSA is merely an artifact of the hydrodynamic model. We address this issue using a different particle-based model below.
Figure 8: Phase diagram of the quorum-sensing ABP model in (a) the BMR with $\lambda=0.5$, $D_{\mathrm{c}}=10$, $L=400$, $t=4\times 10^{2}$ and (b) the AMR with $n_{0}=20$, $\zeta=22$, $L=400$, $t=1.6\times 10^{3}$. The LSA predictions for the phase boundaries (at $\kappa=4.791667$) are indicated by the dotted (H–S boundary), the dash-dotted (H–O boundary), and the dashed (S–O boundary) lines. For comparison, snapshots of particle distribution (with darker regions indicating higher density) obtained by particle-based simulations are shown together. The central coordinates of each snapshot correspond to the parameters used in the simulation. Note that the snapshots in (b) have been cropped to the size of $200$ times $200$ for better visibility.
IV.2 Quorum-sensing ABPs
@@
As a particle-based model more easily comparable with the hydrodynamic counterpart, here we introduce the quorum-sensing ABPs. While their dynamics share Eqs. (2)–(4) with the mechanically interacting ABPs, Eq. (II) is replaced with
$\dot{\mathbf{r}}_{k}(t)=\mu v\big{(}\varrho_{\mathrm{loc}}(\mathbf{r}_{k},t),n(\mathbf{r}_{k},t)\big{)}\mathbf{e}_{k}+\sqrt{2\mu T}\,\bm{\xi}_{k}\;,$ (16)
@@
$v(\varrho_{\mathrm{loc}},n)=\alpha n-\varrho_{\mathrm{loc}}\zeta$ (17)
@@
The kernel $K(\mathbf{r})$ should be positive only within a finite range, which is chosen to be
@@
The quorum-sensing ABPs have some advantages compared to the mechanically interacting ABPs. First, the parameters of the model are identical to those of the hydrodynamic model, making quantitative comparison between the models easier. The only hydrodynamic parameter not explicitly built into the model is $\kappa$, but applying the coarse-graining scheme of Bialké et al. (2013), linearizing the equations with respect to the homogeneous profile, and neglecting higher-order derivatives, we can identify

### v2

@@ Results
We observe that the phase boundary changes discontinuously from $D_{\text{vanilla}}$ to $D_{\text{BMR}}$ as the chemical consumption is turned on. When the chemicals are irrelevant ($\lambda=0$), MIPS requires that $D_{\mathrm{eff}}$ is negative, as discussed above. In contrast, when $\lambda>0$, high local particle density and low local chemical concentration can support each other via positive feedback, which can amplify density and concentration modulations even as diffusive currents governed by positive $D_{\mathrm{eff}}$ try to suppress them. For this reason, the mechanism of MIPS for $\lambda>0$ (which may happen even for $D_{\mathrm{eff}}>0$) is fundamentally different from the corresponding mechanism for $\lambda=0$ (which requires $D_{\mathrm{eff}}<0$), which explains the discontinuous change of the phase boundary.
@@
To check the validity of the LSA, we perform numerical simulations of Eq. (Hydrodynamic model) via the pseudo-spectral method. In Fig. 3(a), we show the phase boundary (indicated by a dotted curve) predicted by the LSA together with snapshots of the numerical simulations at $t=2\times 10^{3}$ as the global mean concentration $n_{0}$ and the deceleration parameter $\zeta$ are varied within the physical regime (the grayed-out area corresponds to the unphysical regime where $v(\rho,n)<0$). The two control parameters are chosen to represent the two major factors contributing to MIPS, namely the self-propulsion $\alpha n_{0}$ and the deceleration $\rho_{0}\zeta$. Note that the coordinates of the center of each snapshot indicate the associated values of $n_{0}$ and $\zeta$. Clearly, the predicted boundary correctly separates the homogeneous snapshot from the phase-separated ones.
We also note that, unlike the vanilla MIPS, chemokinetic effects allow the particles to form clusters (for high $n_{0}$) or bubbles (for low $n_{0}$) even at the same value of $\rho_{0}$. This phenomenon, along with demonstrations of coarsening dynamics, is presented in more detail in Fig. A3, Supplementary Movies 5A, 5B, and Appendix D.
@@
Figure 4: Phase separation behaviors of the hydrodynamic model in the Active Metabolic Regime. Particle density $\rho(\mathbf{r},t)$ (blue) and the square magnitude of local vorticity $|\boldsymbol{\omega}(\mathbf{r},t)|^{2}$ (green), rescaled by its standard deviation, are shown for (a) diffusive droplets ($\lambda=0.0025$, $D_{\mathrm{c}}=1.75$, $n_{0}=10$, $\zeta=22$), (b) ballistic bands ($\lambda=0.01$, $D_{\mathrm{c}}=1$, $n_{0}=17.5$, $\zeta=28.5$), and (c) radiating ripples ($\lambda=0.01$, $D_{\mathrm{c}}=1$, $n_{0}=23.5$, $\zeta=28.5$). Darker green color corresponds to the region with high local vorticity. Through (a–c), particle velocity fields are indicated by red arrows in the insets. (d, e) Mean cluster size in the S phase as the chemical consumption rate $\lambda$, the chemical diffusion coefficient $D_{\mathrm{c}}$, and the system size $L$ are varied. The error bars indicate the standard deviation of cluster size. (f) Mean vorticity as $D_{\mathrm{c}}$ is varied for $\lambda=0.0025$ and (inset) the same quantity as $\lambda$ is varied for $D_{\mathrm{c}}=0.25$. Note that $n_{0}=20$ and $\zeta=22$ in (d–f).
AMR induces small, oscillating clusters
@@
The LSA for the AMR can be done similarly, but it differs from that for the BMR in that $\mathcal{B}$ and $\mathcal{C}$ in Eq. (9) may have different signs ($\mathcal{B}<0$ and $\mathcal{C}>0$). The eigenvalues $\omega_{\pm}$, obtained in Eq. (10), may then have nonzero imaginary parts. Thus, the AMR exhibits the O phase in addition to the H and the S phases, which leads to some extra complications in predicting the phase boundaries. Nonetheless, the LSA can be carried out exactly, as detailed in Appendix B. The analysis yields a set of inequalities determining the range of $D_{\mathrm{eff}}$ belonging to the H phase, which is always broadened by the presence of the particle–chemical coupling ($\lambda>0$) always broadens the H phase. This confirms our previous expectation that the AMR suppresses MIPS.
@@
Examples of the $\zeta$$n_{0}$ and the $D_{\mathrm{c}}$$\lambda$ phase diagrams for the AMR are shown in Fig. 3(b) and (c), respectively. The former parameter set is chosen for direct comparison with Fig. 3(a), and the latter parameter set is chosen to represent how the strength of the particle–chemical coupling $\lambda$ and the time scale of the chemical dynamics $1/D_{\mathrm{c}}$ modify the phase separation behaviors. The LSA prediction for the H–S phase boundary is marked by the dotted line, the H–O boundary by the dash-dotted line, and the S–O boundary by the dashed line. For comparison, we also show the snapshots of the particle density obtained by numerical simulations of Eq. (Hydrodynamic model) via the pseudo-spectral method, where the coordinates of the center of each square corresponds to the parameters used in the simulation. The predicted H–S and the H–O boundaries, for which we expect the LSA to be exact, are in good agreement with the numerics. In contrast, the O phase seems to be broader than indicated by the predicted S–O boundary, as exemplified by the lower left corner of Fig. 3(c). Since the LSA is valid only up to linear order in density and concentration fluctuations, it is only natural that the true S–O boundary would deviate from the LSA prediction.
@@
appearing in Eq. (Hydrodynamic model). We note that high vorticity tends to coincide with regions where patterns evolve more rapidly. With these in mind, we can identify at last three distinct patterns, namely diffusive droplets [see Fig. 4(a)] well within the S phase, ballistic bands near the boundaries of the O phase [see Fig. 4(b)], and radiating ripples well within the O phase [see Fig. 4(c)]. It seems that the latter two patterns are essentially the same, the only difference being that radiating ripples have a smaller length scale. This is also corroborated by the LSA, which shows that the unstable oscillating fluctuations tend to exhibit shorter wavelengths in the regimes of radiating ripples compared to ballistic bands.
@@
Let us first look into the properties of the diffusive droplets. As indicated by Fig. 4(d), the average number of particles forming each droplet (which we call the mean cluster size) decreases with $\lambda$ and increases with $D_{\mathrm{c}}$. The former is in agreement with our intuition that the particle–chemical coupling in the AMR suppresses MIPS. The latter is due to the high $D_{\mathrm{c}}$ strongly suppressing the short-wavelength undulations of the chemical concentration. Since all patterns are formed by the particle density and the chemical concentration sharing the same characteristic length scale, this effect also drives the particles to form larger droplets. However, as implied by the snapshots in Fig. 3(c), the mean cluster size starts to increase again when $D_{\mathrm{c}}$ decreases below $2.5$. This is due to the droplets becoming increasingly motile and actively colliding with each other as the system approaches the S–O boundary.
All these behaviors are in stark contrast to the case of the BMR, where MIPS eventually forms a single macroscopic particle cluster via coarsening. Indeed, Fig. 4(e) indicates that the mean cluster size does not change appreciably with the system size $L$ when $\lambda>0$. In other words, the size of each droplet stays finite, and the S phase of the AMR exhibits microphase separation. This is also supported by the disintegration of a large particle cluster in the S phase, which is further discussed in Fig. A4, Supplementary Movie 6, and Appendix D.
Now, we examine how the system changes quantitatively as it transitions from the S phase with diffusive droplets to the O phase with ballistic bands and radiating ripples. As already shown in Figs. 4(a)–(c), the latter two patterns exhibit spatially extended regions of high vorticity, reflecting their large-scale evolution in time. In contrast, the high-vorticity regions for the diffusive droplets are limited to the droplet boundaries. This implies that the mean vorticity
@@
can be used as an indicator of the O phase. Indeed, Fig. 4(f) shows that the mean vorticity dramatically increases as $D_{\mathrm{c}}$ is decreased for $\lambda=0.0025$, crossing the S–O boundary shown in Fig. 3(c). Notably, even before reaching the S–O boundary, the mean vorticity gradually increases as $D_{\mathrm{c}}$ is decreased, which is then followed by a sharp dip around $D_{\mathrm{c}}=2$. This nonmonotonic behavior might be due to the individual droplets becoming increasingly motile as $D_{\mathrm{c}}$ is reduced, which then merge together by collisions and shorten the boundaries, thus reducing the high-vorticity regions as well.
@@
Figure 5: Profiles of patterns formed by the hydrodynamic model in the Active Metabolic Regime. Particle density $\rho(\mathbf{r})$ and chemical concentration $n(\mathbf{r})$ are shown for (a, b) the S phase with $D_{\mathrm{c}}=2.25,\lambda=0.0025,n_{0}=20,\zeta=22$ and (c, d) the O phase with $D_{\mathrm{c}}=1,\lambda=0.01,n_{0}=25.5,\zeta=17.5$. In (a) and (c), $\rho(\mathbf{r})$ is indicated by heatmaps (blue shades) $n(\mathbf{r})$ by contours, so that the values of $n(\mathbf{r})$ at adjacent contour lines differ by $0.8$. Note that (b) and (d) show cross sections of (a) and (c), respectively. The arrows in (c) and (d) indicate the directions in which the patterns move. See Supplementary Movie 3 for an animated version.
The inset of Fig. 4(f), which shows the mean vorticity for $D_{\mathrm{c}}=0.25$ as $\lambda$ is varied, similarly captures the O phase lying between the S and the H phases in the lower part of Fig. 3(c). Notably, in apparent contrast to Fig. 4(f), the mean vorticity keeps increasing as $\lambda$ approaches the phase boundary. This is because the boundary crossed here is between the O and the H phases. The oscillating patterns, driven by the particle–chemical coupling, move more vigorously with increasing $\lambda$. However, when $\lambda$ increases beyond the H–O boundary, the patterns become transient and quickly die out. For this reason, the mean vorticity exhibits a sharp drop at the H–O boundary.
@@
With the qualitative differences between different patterns thus established, we look into how the particle–chemical coupling creates the traveling patterns in the O phase. When $D_{\mathrm{c}}$ is large, the chemical concentration is slaved to the particle density, so that their undulations overlap almost exactly as shown in Fig. 5(a) and (b). This leads to the diffusive droplets observed in the S phase. In contrast, when $D_{\mathrm{c}}$ is sufficiently small, the particle and the chemical dynamics may have comparable time scales, in which case the following scenario becomes possible. Suppose that the particle density undulations are slightly shifted to the right with respect to those of the chemical concentration. Then, for each density peak, the chemical concentration is lower (higher) on the right (left), decreasing (increasing) the speed of particles there. This means that particles tend to accumulate on the right and evaporate on the left, making the peak travel to the right. The concentration peak follows the density peak but cannot catch up for a prolonged period of time since $D_{\mathrm{c}}$ is small. This is precisely what happens for the density and the concentration fields shown in Figs. 5(c) and (d), which exhibit band-like patterns traveling to the right.
We stress that these traveling patterns in the O phase are consequences of the nonreciprocity reflected in the different signs of $\mathcal{B}$ and $\mathcal{C}$ appearing in Eq. (9). This is similar to how various multispecies active fluids with nonreciprocal interspecies interactions generate traveling patterns [61, 62, 63, 64, 65, 66].
@@
The microphase separation in the AMR is consistent with the behaviors of the particle-based model. However, the distinction between the S and O phases of the AMR was not so pronounced in the mechanically interacting particle-based model. This may be because the microscopic noises, which are neglected in the hydrodynamic model, overwhelm the oscillating patterns in the particle-based model. This raises the question of whether the O phase of the AMR identified by the LSA is merely an artifact of the hydrodynamic model. Below, we address this issue using a different particle-based model, which is expected to have much lower noise than the model with WCA interactions.
Particle-based simulation of quorum-sensing ABPs
@@
As a particle-based model more easily comparable with the hydrodynamic counterpart, here we introduce the quorum-sensing ABPs. While their dynamics share Eqs. (2)–(4) with the mechanically interacting ABPs, Eq. (Particle-based model) is replaced with
$\dot{\mathbf{r}}_{k}(t)=v\big{(}\varrho_{\mathrm{loc}}(\mathbf{r}_{k},t),n(\mathbf{r}_{k},t)\big{)}\mathbf{e}_{k}+\sqrt{2\mu T}\,\bm{\xi}_{k}\;,$ (16)
@@
[…]
@@
Note that, in Eq. (17), $v(\varrho_{\text{loc}})$ is explicitly forbidden from becoming negative. This is because the quorum-sensing ABPs lack short-range repulsion and can form a high-density region where $\alpha n-\varrho_{\mathrm{loc}}\zeta<0$. Meanwhile, the kernel $K(\mathbf{r})$ should be positive only within a finite range, which is chosen to be
@@
Figure 6: Snapshots and phase diagrams of the quorum-sensing active Brownian particles. (a) The Basal Metabolic Regime as the global chemical concentration $n_{0}$ and the deceleration parameter $\zeta$ are varied for $\lambda=0.5$, $D_{\mathrm{c}}=10$, $L=400$, $t=4\times 10^{2}$ and (b) the Active Metabolic Regime as the chemical consumption rate $\lambda$ and the chemical diffusion coefficient $D_{\mathrm{c}}$ are varied for $n_{0}=20$, $\zeta=22$, $L=400$, $t=1.6\times 10^{3}$. The global particle density is $\rho_{0}=1/2$ for both (a) and (b). The unphysical region with $v_{\mathrm{eff}}=\alpha n_{0}-\rho_{0}\zeta<0$ is grayed out. Predictions of linear stability analysis for the phase boundaries (at $\kappa=4.791667$) are indicated by the dotted (H–S boundary), the dash-dotted (H–O boundary), and the dashed (S–O boundary) lines. For comparison, snapshots of particle distribution (with darker regions indicating higher density) obtained by particle-based simulations are shown together. The central coordinates of each snapshot correspond to the parameters used in the simulation. Note that the snapshots in (b) have been cropped to the size of $200\times 200$ for better visibility. See Supplementary Movies 4A and 4B for animated versions of (a) and (b), respectively.
Quorum-sensing ABPs have some advantages over mechanically interacting ABPs. First, the model’s parameters are identical to those of the hydrodynamic model, making their quantitative comparison easier. The only hydrodynamic parameter not explicitly built into the model is $\kappa$, which is an independent parameter in the original hydrodynamic model defined by Eq. (Hydrodynamic model). If we derive yet another hydrodynamic theory for the quorum-sensing ABPs by applying the coarse-graining scheme discussed in Appendix A, the effective value of $\kappa$ for the quorum-sensing model can be expressed in terms of other parameters as follows: